# Ergodicity, KAM, FPUT

### v1

@@ Appendix C FPUT symmetry
[…]
@@
with $(a_{k},\varphi_{k})_{k=1}^{n}$ pairs of conjugate symplectic coordinates $({\bf a},{\mbox{\boldmath$\varphi$}})$.
@@
It is interesting to remark that the the potential of fput-$\beta$-chain with fixed extremes, setting ${\mbox{\boldmath$\sigma$}}=(\sigma_{i})_{i=1}^{4},\sigma_{i}=\pm 1$ and $T_{\bf h}=(\frac{2}{N})^{2}\sum_{k=1}^{\frac{N}{2}}\prod_{j=1}^{4}((1-c_{h_{i}})s_{h_{i}k}-s_{h_{i}}c_{h_{i}k})$ for ${\bf h}=(h_{i})_{i=1}^{4}$, where $c_{h}=\cos\varphi_{h},s_{h}=\sin\varphi_{h}$. can be written:
$V_{4}({\bf Q})=\frac{1}{4}\sum_{{\bf h},{\mbox{\boldmath$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{{a_{h_{j}}}}{{\omega_{h_{j}}}})^{\frac{1}{2}}$ (c.5)
@@
and easily identifies the terms which need to be absent to obtain a normal form which is integrable: if $n_{{\bf h},{\mbox{\boldmath$\sigma$}}}(k)=\sum_{h_{j}=k}\sigma_{j}$ and ${\bf n}=(n_{{\bf h},{\mbox{\boldmath$\sigma$}}}(k))_{k=1}^{N=1}$ the uneliminable resonances $\nu$ are the ones for which $\xi={\bf n}\cdot{\mbox{\boldmath$\omega$}}=0$. Hence to prove the result it is necessary to show that the sum of the terms labeled by the above $\bf n$} for which $\xi=0$ either do not arise (but they do in interesting cases) or that they add up to 0 as a consequence of the symmetry or because ${\bf n}=0$, as found in [18, 28].
Appendix D Toda and FPUT: heuristics
$H({\bf p},{\bf q})=\sum_{k=1}^{N}\Big{(}\frac{1}{2}p_{k}^{2}+\alpha^{2}e^{q_{k}-q_{k+1}}\Big{)}=\sum_{k=1}^{N}(\frac{1}{2}p_{k}^{2}+\alpha^{2}V_{T}(q_{k}-q_{k+1}))$ (d.1)
The periodic Toda lattice model is studied in [31] where $2N$ global canonical action-angle variables $({\bf a},{\mbox{\boldmath$\varphi$}})\in(R_{+}^{N},[0,2\pi]^{N})$ are constructed and evolve at time $t$ into $({\bf a},{\mbox{\boldmath$\varphi$}}+{\mbox{\boldmath$\omega$}}({\bf a})t)$, following a Hamiltonian $H_{\alpha}({\bf a})$. Near ${\bf a}=0$ it is $\omega_{i}({\bf a})=2\alpha\omega_{0,i}+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$, if ${\mbox{\boldmath$\omega$}}_{0}$ is the spectrum of the linear oscillations $\omega_{0,i}=2\sin\frac{\pi i}{N}$.
The $H_{\alpha}({\bf a})$, [31, th.1.4], is globally defined for ${\bf a}\in R^{N}_{+}$ and the Jacobian $J_{\alpha}({\bf a})=\det\frac{{\partial}{\mbox{\boldmath$\omega$}}({\bf a})}{{\partial}{\bf a}}$ is analytic in $\alpha>0$; for $\alpha>0$ it is strictly convex in ${\bf a}\in R_{+}^{N}$ and $J_{\alpha}({\bf a})_{ij}=\frac{1}{4N}\delta_{ij}+O(|{\bf a}|)\neq 0$ and the spectrum is $\omega_{i}=2\alpha\sin(\frac{i\pi}{N})+\frac{1}{4N}a_{i}+O(|{\bf a}|^{2})$.
However application to the fput-chains requires more analysis mainly because the use of the change of variables leads to the Hamiltonian in Eq.c.1 with $\alpha$ small and compare it with Eq.3.6 $\varepsilon$ small: aside from the time rescaling, this means to compare motions with $\alpha$ in Eq.c.1 to motions with $\varepsilon$ in Eq.3.6 setting $\alpha=\varepsilon^{-2}$; so when $\alpha$ is very large.
The analysis in [31] is global and the coordinates $({\bf a},{\mbox{\boldmath$\varphi$}})$ exist for all $\alpha>0$ and transform analytically the Hamiltonian Eq.c.1 into a function $\widetilde{H}_{\alpha}({\bf a})$ with convex Jacobian $J_{ij}=\frac{{\partial}^{2}_{a_{i}}H({\bf a})}{{\partial}^{2}a_{j}}$, see th.(1.1),th.(1.2) in [31]. Hence in the region of couplings relevant for the fput-chains, i.e. $\varepsilon$ close to 0, $\alpha=\varepsilon^{-2}$ large, the Toda lattice Hamiltonian admits action-angles coordinates.
Proceeding heuristically: the Toda lattice can be considered a perturbation of order $\varepsilon^{2}$ of the fput-$\alpha$-lattice, see Eq.3.6:
$H_{3}({\bf p},{\bf q})=\sum_{j=0}^{N-1}\Big{(}\frac{1}{2}(p_{k}^{2}+(q_{k}-q_{k+1})^{2})+\frac{1}{3!}\varepsilon(q_{k}-q_{k+1})^{3}\Big{)}$ (d.2)
Hence the elastic chain perturbed by $h_{3}=\frac{1}{3!}\varepsilon\sum_{j=0}^{N-1}(q_{k}-q_{k+1})^{3}$ and the Toda lattice truncated to order $\varepsilon^{2}$ are the same: thus their first order Birkhoff normal forms, to remove the resonant terms present in $h_{3}$, coincide up to the next order $\varepsilon^{2}$ (the first step in the construction of Birkhoff normal form only involves the common part $h_{3}$). The integrability of the Toda lattice should then imply that the normal form for the above truncation is integrable implying that the normal form of the fput-$\alpha$-chain is integrable.
Therefore the above argument of [26] can be applied (heuristically here, and avoiding any explict calculation of the Birkhoff normal form) to reach the same conclusion and the periodic fput-$\alpha$-chain model, and consequently (again by the symmetry in [26]) the fixed extremes fput-$\beta$-chain, show quasi periodic motions, filling large regions of phase space in any region $\Omega$ where $e<H<E$, up to a set of volume approching 0 as $\varepsilon\to 0$.
The $\alpha$-chain can be studied in the same way, still heuristically: the periodic fput-chain with potential $V_{3,4}(x)=\frac{\varepsilon}{3!}x^{3}+\frac{\varepsilon^{2}}{4!}x^{4}$ is the same as the Toda lattice truncated at second order. Then the Birkhoff’s normal form of second order should be the same as that of the Toda lattice: therefore should be integrable. However the second order normal form of the $V_{3,4}$ fput-chain contains no cross terms between the two contributions to the potential (which have at least order $3$). Hence the two contributions to the mormal form of $V_{3,4}$ fput-chain are the same as those for the simple $V_{3}$-chain ($\alpha$-chain) to second order plus the first order ones of the $V_{4}$ (the $\beta$ chain which is already known to be integrable); hence the integrability of the Toda lattice allowing to express its normal form in terms of the action variables implies the same properties for the Birkhoff’s normal forms of the $V_{3,4}$ and $V_{3}$ fput chains.
The argument can be extended to chains with potential $V(x)=\sum_{k=3}^{s}c_{s}x^{s}$ for $s>4$ but only if $c_{s}=\frac{1}{s!}$, while for $s=3,4$ the $c_{3},c_{4}$ could be quite arbitrary.

### v2

@@ Appendix C FPUT symmetry
$S\Big(\hbox{}\,\vbox{\halign{\strut\hfil$\displaystyle{#}$&$\displaystyle{{}#}$\hfil\cr&{\bf Q_{+}},{\bf Q_{-}}\cr&{\bf P_{+}},{\bf P_{-}}\cr}}\,\Big)=\Big(\hbox{}\,\vbox{\halign{\strut\hfil$\displaystyle{#}$&$\displaystyle{{}#}$\hfil\cr&{\bf Q_{+}},-{\bf Q_{-}}\cr&{\bf P_{+}},-{\bf P_{-}}\cr}}\,\Big),\qquad R\Big(\hbox{}\,\vbox{\halign{\strut\hfil$\displaystyle{#}$&$\displaystyle{{}#}$\hfil\cr&{\bf q}\cr&{\bf p}\cr}}\,\Big)=\Big(\hbox{}\,\vbox{\halign{\strut\hfil$\displaystyle{#}$&$\displaystyle{{}#}$\hfil\cr&{\bf q}^{t}\cr&{\bf p}^{t}\cr}}\,\Big)$ (c.3)
@@
with $(a_{k},\varphi_{k})_{k=1}^{n}$ pairs of conjugate symplectic coordinates $({\bf a},{\mbox{$\varphi$}})$.
@@
It is interesting to remark that the the potential of fput-$\beta$-chain with fixed extremes, setting ${\mbox{$\sigma$}}=(\sigma_{i})_{i=1}^{4},\sigma_{i}=\pm 1$ and $T_{\bf h}=(\frac{2}{N})^{2}\sum_{k=1}^{\frac{N}{2}}\prod_{j=1}^{4}((1-c_{h_{i}})s_{h_{i}k}-s_{h_{i}}c_{h_{i}k})$ for ${\bf h}=(h_{i})_{i=1}^{4}$, where $c_{h}=\cos\varphi_{h},s_{h}=\sin\varphi_{h}$. can be written:
$V_{4}({\bf Q})=\frac{1}{4}\sum_{{\bf h},{\mbox{$\sigma$}}}T_{\bf h}\prod_{j=1}^{4}\sigma_{j}e^{i\sigma_{j}\varphi_{h_{j}}}(\frac{{a_{h_{j}}}}{{\omega_{h_{j}}}})^{\frac{1}{2}}$ (c.5)
@@
and easily identifies the terms which need to be absent to obtain a normal form which is integrable: if $n_{{\bf h},{\mbox{$\sigma$}}}(k)=\sum_{h_{j}=k}\sigma_{j}$ and ${\bf n}=(n_{{\bf h},{\mbox{$\sigma$}}}(k))_{k=1}^{N=1}$ the uneliminable resonances $\nu$ are the ones for which $\xi={\bf n}\cdot{\mbox{$\omega$}}=0$. Hence to prove the result it is necessary to show that the sum of the terms labeled by the above $\bf n$} for which $\xi=0$ either do not arise (but they do in interesting cases) or that they add up to 0 as a consequence of the symmetry or because ${\bf n}=0$, as found in [18, 28].
STATEMENT: This paper has no supplementary data